OpenStudy (anonymous):

You are decorating a cylindrical hat box to give to your grandmother as a gift. You want to paste family photos onto the curved part of the box. The height of the box is 8 inches and the radius is 7 inches. If none of the photos overlap, what is the area of photographs you need? Use three point one four for pi. one hundred seventy-five point eight four square inches three hundred fifty-one point six eight square inches one thousand two hundred thirty point eight eight square inches one thousand four hundred six point seven two square inches

OpenStudy (wolf1728):

Area of a Cylinder's Side = 2* PI * r * height = 2 * 3.14 * 3.5 * 8 = 175.84 square inches The answer is A